OpenStudy (anonymous):

need help with this polynomial function, Please help I need to finish this assignment quick

OpenStudy (anonymous):

Polynomial Function Question 2. Heinz has a list of possible functions. Pick one of the g(x) functions below, show how to find the zeros, and then describe to Heinz the other key features of g(x). o g(x) = x3 – x2 – 4x + 4 o g(x) = x3 + 2x2 – 9x – 18 o g(x) = x3 – 3x2 – 4x + 12 o g(x) = x3 + 2x2 – 25x – 50 o g(x) = 2x3 + 14x2 – 2x – 14

OpenStudy (anonymous):

can some please help me

OpenStudy (anonymous):

@JoshR1213 can you help me

OpenStudy (anonymous):

What's the problem?

OpenStudy (anonymous):

im going to try and find the zeros can u tell me if there right

OpenStudy (anonymous):

If the numbers after the x's are the exponents then you know there has to be 3 zeroes for the first one and since there 1 sign changes when the signs are in their original state and 1 time when you change them you know there are 1 positive and 1 negative zeroes and a complex

OpenStudy (anonymous):

If i did that right, this wasn't my strongest subject

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

well I finished anyways thanks